OpenStudy (anonymous):

what is the correct expression to: 2cos2(θ)−1 =?

OpenStudy (anonymous):

\[2\cos ^{2}(\theta)-1 ....... or ........ 2\cos(2\theta)-1\] Which form?

OpenStudy (anonymous):

Enter the expression 2cos2(θ)−1, where θ is the lowercase Greek letter theta.

OpenStudy (anonymous):

the 1st form you wrote

OpenStudy (anonymous):

What exactly is asked? Is the expression equal to something?

OpenStudy (cwrw238):

cos 2θ = 2cos^2θ - 1

OpenStudy (cwrw238):

the question marks are theta